OpenStudy (anonymous):

Find a formula for the upper sum by dividing the interval into n equal subintervals. Then take a limit of these sums as n goes to infinity to calculate the area under the curve: f(x)=3x^2 +x over the interval [0,1] sum_{i=1}^{n} i= n(n+1)/2, sum_{i=1}^{n}= n(n+1)(2n+1)/6

OpenStudy (anonymous):

\[\sum_{i=1}^{n} i= n(n+1)/2, \sum_{i=1}^{n}= n(n+1)(2n+1)/6\sum\]

OpenStudy (anonymous):

\[\sum_{i=1}^{n} i= n(n+1)/2, \sum_{i=1}^{n}= n(n+1)(2n+1)/6\], this is the correct form

OpenStudy (zarkon):

\[\sum_{i=1}^{n}i^2=\frac{n(n+1)(2n+1)}{6}\]

OpenStudy (anonymous):

when i tried typing that, it didn't look as nice as yours, lol

OpenStudy (zarkon):

\sum_{i=1}^{n}i^2=\frac{n(n+1)(2n+1)}{6}

OpenStudy (anonymous):

oh okay